Problem 1 (10 marks)

Suppose GDP equals $300 trillion, consumption equals $24 trillion, the government spends $3 trillion and has a budget deficit of $500 billion.

  • Find public saving, taxes, private saving, national saving, and investment.
Answer #1

Answer

Taxes =govenmet spending -budget deficit

=3-0.5

=2.5 trillion

======

Pubcic saving =T-G=2.5-3=-0.5=-500 billion

===

Investment =Y-C-G=natioanl saving

=300-24-3

=273 trillion

====

Private saving =Y-T-C

=300-2.5-24

=273.5 trilion